This BubbleSort algorithm takes as its input the list of words w), W2, ...,W.. Step 1: Let j=1. Step 2: If j=n then output
Answer #1

This BubbleSort algorithm takes as its input the list of words w), W2, ...,W.. Step 1: Let j=1. Step 2: If j=n then output

2. This list Bubble sort algorithm takes are its input the of words W, W2, ... Who - Anso for the Bubble sort algorithm, giso total number of comparisons will be =(n-1)+(n-2)+6–3) +--... +(4)+(3) + (2)+(1) = n(n-1)/2 • Number of So total subtractioa Step 6 - Replace j by jt 1 Step 7% Go to step 2 the total number from the This function will then return of comparisons and